OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 5 cos 3θ

OpenStudy (anonymous):

is there a pic?

OpenStudy (sidsiddhartha):

it will be a rose with 3 petals http://www.wolframalpha.com/input/?i=r%3D5cos3θ&lk=4

OpenStudy (anonymous):

ok, then it will be the x-axis

OpenStudy (sidsiddhartha):

to check symmetry about y repalce theta by (pi-theta) so r=5cos3θ θ=(π−θ) r=5cos(3(π−θ)) r=5cos(3π−3θ) so it does not resembles original equation

OpenStudy (sidsiddhartha):

test for x-symmetry by setting θ=−θ